### v1

@@
Dust in galaxies is an important tracer of galaxy properties and their evolution over time. The physical origin of the grain size distribution, the dust chemical composition and, hence, the associated ultraviolet-to-optical extinctions in diverse galaxies remains elusive. To address this issue, we introduce a model for dust evolution in the ramses code for simulations of galaxies with a resolved multiphase interstellar medium. Dust is modelled as a fluid transported with the gas component, and is decomposed into two sizes, 5 nm and $0.1\,\mu\rm m$, and two chemical compositions for carbonaceous and silicate grains. This dust model includes the growth of dust by accretion of elements from the gas phase and by the release of dust in stellar ejecta, the destruction by thermal sputtering, supernovae, and astration, and the exchange of dust mass between the two main populations of grain sizes by coagulation and shattering. Using a suite of isolated disc simulations with different masses and metallicities, the simulations can explore the role of these processes in shaping the key properties of dust in galaxies. The simulated Milky Way analogue reproduces the dust-to-metal mass ratio, depletion factors, size distribution and extinction curves of the Milky Way. Galaxies with lower metallicities reproduce the observed decrease in the dust-to-metal mass ratio with metallicity at around a few $0.1\,\rm Z_{\odot}$. This break in the DTM corresponds to a galactic gas metallicity threshold that marks the transition from an ejecta-dominated to an accretion-dominated grain growth, and that is different for silicate and carbonaceous grains, with $\simeq 0.1\,\rm Z_{\odot}$ and $\simeq 0.5\,\rm Z_{\odot}$ respectively. This leads to more Magellanic Cloud-like extinction curves, i.e. with steeper slopes in the ultraviolet and a weaker bump feature at $2175\,\AA$, in galaxies with lower masses and lower metallicities. Steeper slopes in these galaxies are caused by the combination of the higher efficiency of gas accretion by silicate relative to carbonaceous grains, and by the low rates of coagulation that preserves the amount of small silicate grains. Weak bumps are due to the overall inefficient accretion growth of carbonaceous dust at low metallicity, whose growth is mostly supported by the release of large grains in SN ejecta. We also show that the formation of CO molecules is a key component to limit the ability of carbonaceous dust to grow, in particular in low-metallicity gas-rich galaxies.
@@
Yohan Dubois 11    Francisco Rodríguez Montero 22    Corentin Guerra 33    Maxime Trebitsch 44    San Han 55
Ricarda Beckmann
@@
   Sukyoung K. Yi 55    Joseph Lewis 11    and J. K. Jang 55
(Received / Accepted )
@@ 1 Introduction
While galaxy attenuation curves offer insights into the dust properties (Calzetti et al. 2000), it is crucial to note that they result from the interplay of complex radiative transfer effects within the intricate geometry of the interstellar medium (ISM) and stellar distribution (e.g., Gordon et al. 1997; Witt & Gordon 2000; Granato et al. 2000; Inoue 2005; Narayanan et al. 2018). A key source of knowledge about dust properties comes from the extinction curves of multiple lines-of-sights of individual stars (for which the complexity of dust absorption is better characterised) in the Milky Way or from the Magellanic Clouds. The Milky Way extinction curve is characterised by a moderate extinction in the ultraviolet, as opposed to Magellanic Clouds, and by a significant feature in the optical range, called the $2175\,\AA$ bump, very faint in the Large Magellanic Cloud and almost indistinguishable in the Small Magellanic Cloud (Pei 1992). Thanks to the description with Mie theory of light absorption and scattering by solid spherical bodies, and optical properties of grains, it is possible to interpret the shape of the Milky Way extinction curve by a number density distribution skewed towards small grains (a few nm in size) but dominated by large grains ($0.1\,\mu\rm m$) in mass, as well as a mixed composition of carbonaceous grains and of silicate grains (Weingartner & Draine 2001a). Although, the exact nature of the carbonaceous grains (amorphous, aromatic, aliphatic) and silicate grains (olivine, pyroxene, iron inclusions) and how they mix up in a single grain body is complex (Jones et al. 2017), it is possible to make predictions about the grain properties as a function of the galaxy properties and test them against observables (e.g. Mathis et al. 1977; Draine & Lee 1984; Weingartner & Draine 2001a; Zubko et al. 2004; Draine & Li 2007; Compiègne et al. 2011; Galliano et al. 2018), such as the extinction or attenuation laws, or the depletion of the various elements in the gas phase.
[…]
@@
Several dust evolution models in galaxies have been developed which include the dust released by stars, the growth of dust by accretion of elements from the gas phase, and the destruction of dust by astration, supernovae and thermal sputtering. In the seminal work of Dwek (1998), they developed a one zone analytical model to follow the evolution of carbonaceous and silicate grains separately, and they have been able to reproduce the observed amount of dust in Milky Way-like galaxies. Such one zone models have been further refined to track the distribution of grains in sizes including the coagulation and shattering of grains (Asano et al. 2013b). Evolution of dust, followed as a separate component to that of metals from the gas phase, is also modelled in complete hydrodynamical simulations of galaxies (e.g. Bekki 2013; McKinnon et al. 2016; Aoyama et al. 2017; Gjergo et al. 2018; Li et al. 2019, 2021; Graziani et al. 2020; Granato et al. 2021; Trebitsch et al. 2021; Choban et al. 2022). Some of these hydrodynamical simulations make various assumptions about the dust physics, either neglecting the chemical differentiation of grains (as in e.g. Aoyama et al. 2017) or the grain sizes (as in e.g. Graziani et al. 2020; Choban et al. 2022), or both (McKinnon et al. 2016; Trebitsch et al. 2021), or using a unique size distribution for different dust grain species (Li et al. 2021), or using too low resolution to separate ab initio the dense gas structures, in which accretion and coagulation efficiently proceed, from the diffuse ISM (Gjergo et al. 2018; Granato et al. 2021; Li et al. 2021).
@@
Our goal in this paper is to provide a theoretical understanding of the physical origin of the dust properties and, hence, of the different shapes of extinction curves observed as in the Milky Way, or the Magellanic Clouds. Therefore, our approach is to implement and test against various observables a model of dust evolution in the adaptive mesh refinement code ramses (Teyssier 2002) that evolves the sizes and the chemical composition of dust. This model will assume a two-size grain approximation, at $5\,\rm nm$ and at $0.1\,\rm\mu m$, following separately the evolution of carbonaceous and silicate grains, and will be tested in numerical simulations with sufficient resolution to capture the structure of a turbulent multiphase ISM.
@@
The paper is structured as follows. In section 2, we provide an overview of the galactic physics (cooling, star formation, supernova feedback, chemical enrichment) used in ramses, and setup of initial conditions to simulate isolated disc galaxies. In section 3, we describe the details of the model of dust evolution. In section 4, we test our model in a typical Milky Way-like galaxy including stringent tests on extinction curves. In section 5, we extend our predictions to various galaxy masses and explain how Magellanic Cloud-like extinction curves can be obtained. We discuss the limitations of this work in section 6 and, in section 7, we outline our conclusions.
@@ 2.1 Initial conditions and resolution
Table 1: Galaxy initial conditions identifier with halo mass $M_{\rm h}$, gas mass $M_{\rm gas}$, disc stellar mass $M_{\rm\star,d}$, bulge stellar mass $M_{\rm\star,b}$, gas fraction in the galaxy $f_{\rm gas}$, baryon fraction in the halo $f_{\rm bar}$, and disc scaling radius $R_{\rm d}$.
@@
Simulations are performed with the adaptive mesh refinement code ramses (Teyssier 2002). It solves the hydrodynamics on a cartesian grid with refinement with the Godunov method using the unsplit MUSCL-Hancock scheme with the minmod slope limiter to preserve the total variation diminishing property of the scheme, and the Harten-Lax-van Leer-Contact approximate Riemann solver to obtain the flux at a cell interface. Dark matter and stars are represented by particles that are evolved with a leap-frog scheme. The gravitational potential is obtained by a particle-mesh method where particles are projected on the grid with a cloud-in-cell interpolation.
@@
The spatial resolution is varied from $\Delta x=72\,\rm pc$ to $9\,\rm pc$ by multiples of 2 with a resolution of $18\,\rm pc$ for our fiducial simulations. The halo mass is either $M_{\rm h}=9\times 10^{9}$, $9\times 10^{10}$, or $9\times 10^{11}\,\rm M_{\odot}$ with a total galaxy mass (gas and stars) that represents 4.4 per cent of the halo virial mass, hence that is respectively $M_{\rm gal}=3.8\times 10^{8}$, $3.8\times 10^{9}$, or $3.8\times 10^{10}\,\rm M_{\odot}$, for simulations that we name respectively G8, G9 and G10. The halo, stellar disc and bulge are initialised with respectively $10^{6}$, $10^{6}$ and $10^{5}$ particles. We allow refinement where the total mass exceeds $8\times 10^{3}\,\rm M_{\odot}$ starting at level 6. In addition to pseudo-Lagrangian mass refinement, we add a refinement criterion on the Jeans length so that it is refined with at least 4 cells down to the maximum level of refinement. The box sizes of respectively 150, 300, and $600\,\rm pc$ for G8, G9, and G10 are evolved with outflowing boundary conditions.
@@
The dark matter (DM), star and gas distributions are initialised with makedisk (Springel et al. 2005), following the strategy introduced in Rosdahl et al. (2017) for ramses. The DM mass distribution follows a Navarro et al. (1997) profile with concentration $c=10$ and a spin parameter of $\lambda_{\rm DM}=0.04$ (Macciò et al. 2008). The disc distribution is decomposed into a stellar and a gaseous component, with density profiles of $\rho\propto e^{-R/R_{\rm d}}{\rm sech}^{2}(z/(2H))$, where $R$ is the cylindrical radius, $R_{\rm d}$ is the disc scale length, $z$ is the disc height, and $H$ is the disc characteristic scale height that we take as $H=0.1R_{\rm d}$. The bulge follows a Hernquist density profile with bulge scale length equals to $H$. The gas disc density profile with a temperature of $10^{4}\,\rm K$ is used up to a cut radius and a height radius of respectively $R_{\rm c}=5R_{\rm d}$ and $H_{\rm c}=5H$ after which, a uniform density of $10^{-6}\,\rm H\,cm^{-3}$ and a uniform temperature of $10^{7}\,\rm K$ are imposed. For the G8, G9, and G10 galaxies, we use respectively a disc scale radius of $0.7$, $1.5$ and $3.2$ kpc.
The gas metallicity is initialised with a decreasing radial profile, following
@@
where $R_{c}=7.42\,\rm kpc$, and $Z_{0}$ is the scaling metallicity of the profile, which takes values of $Z_{0}=0.045,0.15,0.45,0.90\,\rm Z_{\odot}$, with $Z_{\odot}=0.01345$, which give respectively a mean metallicity of the gas of $0.1,0.3,1,3\,\rm Z_{\odot}$. The various elements followed in this work (H, He, C, N, O, Mg, Fe, Si, and S) are initialised in the gas phase with solar mass fractions (Asplund et al. 2009). The exact mass fraction of each individual element is a function of the metallicity and of star formation history of the galaxy, as can be observed for low metallicity stars in the Milky Way. We assume that 10 per cent of the initial refractory [^1] Through an intentional terminological adjustment, we will employ the term ‘refractory’ throughout this text to refer to the atomic elements contributing to the composition of the dust grains in our model, i.e. this includes Mg, Fe, Si, O, and C. elements are trapped in dust grains for the G10LG galaxy and 0.1 per cent for the lower mass galaxies G8 and G9, whose mass is equally distributed in sizes, which corresponds to a ${\rm DTM}=3.6\times 10^{-2}$ and $3.6\times 10^{-4}$ respectively.
For galaxies simulated with the large gas fractions (i.e. galaxy G8HG, G9HG, and G10HG), a special procedure to smoothly relax the initial conditions is required. During the first 10s of Myr, the gas looses pressure support due to rapid cooling and produces a strong burst of star formation rate (SFR). In turn this burst of SFR creates a massive outflow that can remove very significant amounts of gas from the galaxy and turn a gas-rich galaxy into a gas-poor galaxy. In order to circumvent this shortcoming, and after some tests, we have introduced a perturbation in the initial vertical velocity field with an amplitude of $30\,\rm km\,s^{-1}$ modulated by two sinusoids in the $x$ and $y$ axis of the cartesian box (the $z$ axis is oriented along the spin axis of the disc) with wavelength of $200\,\rm pc$. In addition to this initial perturbation of the velocity field, we also impose that the star formation efficiency (see section 2.3) linearly ramps up with time with a time scale of 100 Myr.

### v2

@@
[…]
@@
[…]
(Received 28 February 2024 / Accepted 30 April 2024)
Dust in galaxies is an important tracer of galaxy properties and their evolution over time. The physical origin of the grain size distribution, the dust chemical composition, and, hence, the associated ultraviolet-to-optical extinctions in diverse galaxies remains elusive. To address this issue, we introduce a model for dust evolution in the ramses code for simulations of galaxies with a resolved multiphase interstellar medium. Dust is modelled as a fluid transported with the gas component, and is decomposed into two sizes, 5 nm and $0.1\,\mu\rm m$, and two chemical compositions for carbonaceous and silicate grains. This dust model includes the growth of dust by accretion of elements from the gas phase and by the release of dust in stellar ejecta, the destruction by thermal sputtering, supernovae, and astration, and the exchange of dust mass between the two main populations of grain sizes by coagulation and shattering. Using a suite of isolated disc simulations with different masses and metallicities, the simulations can explore the role of these processes in shaping the key properties of dust in galaxies. The simulated Milky Way analogue reproduces the dust-to-metal mass ratio, depletion factors, size distribution and extinction curves of the Milky Way. Galaxies with lower metallicities reproduce the observed decrease in the dust-to-metal mass ratio with metallicity at around a few $0.1\,\rm Z_{\odot}$. This break in the dust-to-metal ratio corresponds to a galactic gas metallicity threshold that marks the transition from an ejecta-dominated to an accretion-dominated grain growth, and that is different for silicate and carbonaceous grains, with $\simeq 0.1\,\rm Z_{\odot}$ and $\simeq 0.5\,\rm Z_{\odot}$ respectively. This leads to more Magellanic Cloud-like extinction curves, i.e. with steeper slopes in the ultraviolet and a weaker bump feature at $2175\,\AA$, in galaxies with lower masses and lower metallicities. Steeper slopes in these galaxies are caused by the combination of the higher efficiency of gas accretion by silicate relative to carbonaceous grains and by the low rates of coagulation that preserves the amount of small silicate grains. Weak bumps are due to the overall inefficient accretion growth of carbonaceous dust at low metallicity, whose growth is mostly supported by the release of large grains in SN ejecta. We also show that the formation of CO molecules is a key component to limit the ability of carbonaceous dust to grow, in particular in low-metallicity gas-rich galaxies.
@@ 1 Introduction
Unlike extinction, corresponding to a line-of-sight emitted by a single point source whose light is absorbed and scattered, the attenuation is the result of the collection of multiple sources with different extinctions into a single line-of-sight. While galaxy attenuation curves offer insights into the dust properties (Calzetti et al. 2000), it is crucial to note that they result from the interplay of complex radiative transfer effects within the intricate geometry of the interstellar medium (ISM) and stellar distribution (e.g. Gordon et al. 1997; Witt & Gordon 2000; Granato et al. 2000; Inoue 2005; Narayanan et al. 2018). A key source of knowledge about dust properties comes from the extinction curves of multiple lines-of-sights of individual stars (for which the complexity of dust absorption is better characterised) in the Milky Way or from the Magellanic Clouds. The Milky Way extinction curve is characterised by a moderate extinction in the ultraviolet, as opposed to Magellanic Clouds, and by a significant feature in the optical range, called the $2175\,\AA$ bump, very faint in the Large Magellanic Cloud and almost indistinguishable in the Small Magellanic Cloud (Pei 1992). Thanks to the description with Mie theory of light absorption and scattering by solid spherical bodies, and optical properties of grains, it is possible to interpret the shape of the Milky Way extinction curve by a number density distribution skewed towards small grains (a few nm in size) but dominated by large grains ($0.1\,\mu\rm m$) in mass, as well as a mixed composition of carbonaceous grains and of silicate grains (Weingartner & Draine 2001a). Although, the exact nature of the carbonaceous grains (amorphous, aromatic, aliphatic) and silicate grains (olivine, pyroxene, iron inclusions) and how they mix up in a single grain body is complex (Jones et al. 2017), it is possible to make predictions about the grain properties as a function of the galaxy properties and test them against observables (e.g. Mathis et al. 1977; Draine & Lee 1984; Weingartner & Draine 2001a; Zubko et al. 2004; Draine & Li 2007; Compiègne et al. 2011; Galliano et al. 2018; Hensley & Draine 2023), such as the extinction or attenuation laws, or the depletion of the various elements in the gas phase.
Not only is dust key to reconstruct the properties of galaxies from their SED, but it also intervenes directly or indirectly in the thermodynamics of the ISM. Dust is a catalyst for molecular H2 formation, that is an important tracer of star formation with the so-called Kennicutt-Schmidt relation (Kennicutt 1998), and which is a key component for gas cooling by rotational transitions or for heating through H2 formation (Hollenbach & McKee 1979). Also dust directly affects the heating of the gas via photo-electric heating, or the cooling of the gas by depletion of elements in the gas phase, by dust cooling at low temperatures in the dark cores of molecular clouds, or by energy exchange through collisions with electrons in X-ray bright gas. As a contributor to the reprocessing of light in the ISM, dust can have a role in radiative feedback at galactic scales. In particular, the role of dust is certainly exacerbated in massive starburst galaxies which appear as ultra-luminous IR galaxies (or sub-millimetre galaxies at high redshift), as they can obscure up to 99% of their UV-visible emission (Casey et al. 2014).
@@
Several dust evolution models in galaxies have been developed which include the dust released by stars, the growth of dust by accretion of elements from the gas phase, and the destruction of dust by astration, supernovae and thermal sputtering. In the seminal work of Dwek (1998), they developed a one zone analytical model to follow the evolution of carbonaceous and silicate grains separately, and they were able to reproduce the observed amount of dust in Milky Way-like galaxies. Such one zone models were further refined to track the distribution of grains in sizes including the coagulation and shattering of grains (Asano et al. 2013b). Evolution of dust, followed as a separate component to that of metals from the gas phase, is also modelled in complete hydrodynamical simulations of galaxies (e.g. Bekki 2013; McKinnon et al. 2016; Aoyama et al. 2017; Gjergo et al. 2018; Li et al. 2019, 2021; Graziani et al. 2020; Granato et al. 2021; Trebitsch et al. 2021; Choban et al. 2022). Some of these hydrodynamical simulations make various assumptions about the dust physics, either neglecting the chemical differentiation of grains (as in e.g. Aoyama et al. 2017) or the grain sizes (as in e.g. Graziani et al. 2020; Choban et al. 2022), or both (McKinnon et al. 2016; Trebitsch et al. 2021), or using a unique size distribution for different dust grain species (Li et al. 2021), or using too low resolution to separate ab initio the dense gas structures, in which accretion and coagulation efficiently proceed, from the diffuse ISM (Gjergo et al. 2018; Granato et al. 2021; Li et al. 2021).
@@
Our goal in this paper is to provide a theoretical understanding of the physical origin of the dust properties and, hence, of the different shapes of extinction curves observed as in the Milky Way, or the Magellanic Clouds. Therefore, our approach is to implement and test against various observables a model of dust evolution in the adaptive mesh refinement code ramses (Teyssier 2002) that evolves the sizes and the chemical composition of dust. This model assumes a two-size grain approximation, at $5\,\rm nm$ and at $0.1\,\rm\mu m$, following separately the evolution of carbonaceous and silicate grains, and is tested in numerical simulations with sufficient resolution to capture the structure of a turbulent multiphase ISM.
@@
The paper is structured as follows. In section 2, we provide an overview of the galactic physics (cooling, star formation, supernova feedback, chemical enrichment) used in ramses, and the setup of initial conditions to simulate isolated disc galaxies. In section 3, we describe the details of the model of dust evolution. In section 4, we test our model in a typical Milky Way-like galaxy including stringent tests on extinction curves. In section 5, we extend our predictions to various galaxy masses and explain how Magellanic Cloud-like extinction curves can be obtained. We discuss the limitations of this work in section 6 and, in section 7, we outline our conclusions.
@@ 2.1 Initial conditions and resolution
Table 1: Galaxy initial conditions’ identifier and their properties.
@@
Our simulations were performed with the adaptive mesh refinement code ramses (Teyssier 2002). It solves the hydrodynamics on a cartesian grid with refinement with the Godunov method using the unsplit MUSCL-Hancock scheme with the minmod slope limiter to preserve the total variation diminishing property of the scheme, and the Harten-Lax-van Leer-Contact approximate Riemann solver to obtain the flux at a cell interface. Dark matter and stars are represented by particles that are evolved with a leap-frog scheme. The gravitational potential is obtained by a particle-mesh method where particles are projected on the grid with a cloud-in-cell interpolation.
@@
The spatial resolution was varied from $\Delta x=72\,\rm pc$ to $9\,\rm pc$ by multiples of 2 with a resolution of $18\,\rm pc$ for our fiducial simulations. The halo mass was either $M_{\rm h}=10^{10}$, $10^{11}$, or $10^{12}\,\rm M_{\odot}$, with a total galaxy mass (gas and stars) that represented $f_{\rm bar}=3.8\,\%$ of the halo mass, hence that was respectively $M_{\rm gal}=3.8\times 10^{8}$, $3.8\times 10^{9}$, or $3.8\times 10^{10}\,\rm M_{\odot}$, for simulations that we named respectively G8, G9 and G10. The halo, stellar disc and bulge were initialised with respectively $10^{6}$, $10^{6}$ and $10^{5}$ particles. We allowed refinement where the total mass exceeded $8\times 10^{3}\,\rm M_{\odot}$ starting at level 6. In addition to pseudo-Lagrangian mass refinement, we added a refinement criterion on the Jeans length so that it is refined with at least 4 cells down to the maximum level of refinement. The box sizes of respectively 150, 300, and $600\,\rm pc$ for G8, G9, and G10 were evolved with outflowing boundary conditions.
@@
The dark matter (DM), star and gas distributions were initialised with makedisk (Springel et al. 2005), following the strategy introduced in Rosdahl et al. (2017) for ramses. The DM mass distribution followed a Navarro et al. (1997) profile with concentration $c=10$ and a spin parameter of $\lambda_{\rm DM}=0.04$ (Macciò et al. 2008). The disc distribution was decomposed into a stellar and a gaseous component (with mass $M_{\star,\rm d}$ and $M_{\rm gas}$ resp.), with density profiles of $\rho\propto e^{-R/R_{\rm d}}{\rm sech}^{2}(z/(2H))$, where $R$ is the cylindrical radius, $R_{\rm d}$ is the disc scale length, $z$ is the disc height, and $H$ is the disc characteristic scale height that we took as $H=0.1R_{\rm d}$. The stellar bulge component (with mass $M_{\star,\rm b}$) followed a Hernquist density profile with bulge scale length equals to $H$. The gas disc density profile with a temperature of $10^{4}\,\rm K$ was used up to a cut radius and a height radius of respectively $R_{\rm c}=5R_{\rm d}$ and $H_{\rm c}=5H$ after which, a uniform density of $10^{-6}\,\rm H\,cm^{-3}$ and a uniform temperature of $10^{7}\,\rm K$ were imposed. For the G8, G9, and G10 galaxies, we used respectively a disc scale radius of $0.7$, $1.5$, and $3.2$ kpc. We varied the gas fraction $f_{\rm gas}=M_{\rm gas}/(M_{\rm gas}+M_{\star,\rm d}+M_{\star,\rm b})$ between two extreme values. The initial properties of the simulated galaxies and their halos are shown in table 1.
The gas metallicity was initialised with a decreasing radial profile, following
@@
where $R_{c}=7.42\,\rm kpc$, and $Z_{0}$ is the scaling metallicity of the profile, which took values of $Z_{0}=0.045,0.15,0.45,0.90\,\rm Z_{\odot}$, with $Z_{\odot}=0.01345$, which gave respectively a mean metallicity of the gas of $0.1,0.3,1,3\,\rm Z_{\odot}$. The various elements followed in this work (H, He, C, N, O, Mg, Fe, Si, and S) were initialised in the gas phase with solar mass fractions (Asplund et al. 2009). The exact mass fraction of each individual element is a function of the metallicity and of star formation history of the galaxy, as can be observed for low metallicity stars in the Milky Way, but we neglected this aspect. We assumed that 10 per cent of the initial refractory [^1] Through an intentional terminological adjustment, we will employ the term ‘refractory’ throughout this text to refer to the atomic elements contributing to the composition of the dust grains in our model, i.e. this includes Mg, Fe, Si, O, and C. elements were trapped in dust grains for the G10LG galaxy and 0.1 per cent for the lower mass galaxies G8 and G9, whose mass were equally distributed in sizes, which corresponded to a ${\rm DTM}=3.6\times 10^{-2}$ and $3.6\times 10^{-4}$ respectively.
For galaxies simulated with the large gas fractions (i.e. galaxy G8HG, G9HG, and G10HG), a special procedure to smoothly relax the initial conditions was required. Due to the very idealised nature of the disc initial conditions, during the first 10s of Myr, the gas loses pressure support due to rapid cooling and produces a strong burst of star formation rate (SFR). In turn this burst of SFR creates a massive outflow that can remove very significant amounts of gas from the galaxy and turn a gas-rich galaxy into a gas-poor galaxy. In order to circumvent this shortcoming, and after some tests, we introduced a perturbation in the initial vertical velocity field with an amplitude of $30\,\rm km\,s^{-1}$ modulated by two sinusoids in the $x$ and $y$ axis of the cartesian box (the $z$ axis is oriented along the spin axis of the disc) with wavelength of $200\,\rm pc$. In addition to this initial perturbation of the velocity field, we also imposed that the star formation efficiency (see section 2.3) linearly ramped up with time with a timescale of 100 Myr.